# The Preheating Stage on The Starobinsky Inflation after ACT

### v3

@@ 5 On the fermionic preheating
In this part, we will discuss the possibility that, instead of the scalar $\chi$, the scalaron is coupled with fermions. So far, we have only discussed the bosonic preheating, even though fermionic preheating is equally possible. This type of preheating has already been discussed in, e.g., Refs. [greene2000fermionic, greene1999fermion, giudice1999fermion, peloso2000fermion]. In this section, we will qualitatively discuss the possibility of this type of preheating.
If the scalaron is only coupled with certain fermions, at a large coupling [greene1999fermion], the preheating is similar to the bosonic case. However, in our paper, it is heavily suppressed by the Pauli exclusion principle, in contrast to the Bose-Einstein condensate (BEC) that can occur during bosonic preheating. Fermion production becomes less suppressed once the universe expands. However, even after the end of preheating, the potential of the scalaron remains quadratic ($\propto\phi^{2}$), which still favors a matter-dominated regime. Fermions are more likely to be produced at the end of the preheating phase due to the looseness of the Pauli blocking. Finally, we safely said that preheating via this mode is strongly disfavored.
@@ 6 The PBH formation from preheating
In this section, we will discuss PBH formation due to preheating in the Starobinsky model. We will show later that the PBH formation is consistent with it being produced during the early stage of preheating.
@@
By using $\beta_{\text{PBH}}=10^{-20}$ and $\delta=\delta_{c}=0.7$ [green2001primordial, niemeyer1999dynamics] on Eq. (6.1), we obtain the mass variance $\sigma_{\delta}$ as
@@
For the production of PBH during preheating, the strong parametric resonance should be applied. Our assumptions implied that PBH is most likely produced when [liddle2000super, green2001primordial]
@@
The power spectrum of the variable $x$ can be expressed as [liddle2000cosmological]
@@
where $k_{x}=|\textbf{k}_{x}|$ represents the comoving wavenumber associated with $x$. Note that $x_{k}$ denotes the inverse Fourier transform of $x$. The influence of preheating on $\chi$ production can be determined as [kofman1997towards, liddle2000super, green2001primordial]
@@
as the comoving wavenumber evaluated at the Hubble exit at the end of inflation. For $Q\gg 1$, in which it is required to get the largest production, we get $\mu_{p}=\mu_{0}\simeq\frac{1}{2\pi}\ln 3$. In addition, we will calculate the non-adiabatic power spectrum of $\chi$ as [liddle2000super, green2001primordial]
@@
Note that $\vartheta$ is the angle between $p$ and $p^{\prime}$. To solve the last equation, we need to impose [green2001primordial]
@@
Finally, the mass variance $\sigma_{\delta}$ during the matter-dominated era can be calculated as follows [green2004NewCalculation]
@@
Table 1: The duration of the required PBH abundance. In this table, we used $\xi=1$. Based on the table, it seems that the $\Delta t$ is insensitive to the chosen parameters.
Eq. (6.15) shows the dependence on $k_{\chi}$, $\xi$, $m_{\chi}$, and $\Delta t$. In the following, we present the numerical results for Eq. (6.15) in Table 1. From our numerical analysis, by varying $k_{\chi}$ and $m_{\chi}$ with fixed $\xi$, it appears that $\Delta t$ is insensitive to the chosen parameters. The duration required to obtain a sufficient amount of PBHs is found to be $\sim 10^{7}M_{p}^{-1}$. This value is consistent with Fig. 2, which shows that the effective growth of $\chi$ production occurs after $t>10^{6}M_{p}^{-1}$. Note that in our numerical calculation, we keep $m_{\chi},k_{\chi}\ll\xi M^{2}$. In this case, the PBH constraint may not jeopardize our model. Moreover, even if $\chi$ grows larger, at some point it cannot collapse into PBHs due to the rapidly expanding universe. Thus, PBHs are suppressed beyond a certain stage [green2001primordial].
@@ 7.1 The Perturbative decay of the scalaron to scalar $\chi$
$\sqrt{\frac{3}{2}}\frac{\phi}{M_{p}}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi,$ (7.1)
@@
$T_{\text{reh}}\approx 3.14\times 10^{9}\text{GeV}.$ (7.4)
@@ 7.2 The Perturbative decay of the scalaron to fermion $\psi$
$\mathcal{L}_{\text{Dirac}}=e^{-3\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m_{\psi})\psi,$ (7.5)
@@
$3\sqrt{\frac{2}{3}}\frac{m_{\psi}}{M_{p}}\phi\bar{\psi}\psi.$ (7.6)
@@
$\Gamma(\phi\rightarrow\bar{\psi}\psi)\simeq\frac{3}{4\pi}\frac{m_{\psi}^{2}M}{M_{p}^{2}},$ (7.7)
@@
$T_{\text{reh}}\simeq 8.07\times 10^{-4}m_{\psi}.$ (7.8)
By using the ACT reheating temperature constraint of $10^{4}$ GeV, $m_{\psi}\sim 10^{8}$ GeV could satisfy the prediction. Generally speaking, this mechanism is favored but we will see later that at some point, a certain mechanism is more favored.
@@ 7.3 Remarks on the perturbative mode
In our discussion about the perturbative reheating by scalar $\chi$ and fermion $\psi$, it is shown that the scalar mode is disfavored by ACT results as it produces a high reheating temperature. Furthermore, as it is enhanced by BEC, the reheating temperature from this mode could be larger than $10^{9}$ GeV. On the other hand, reheating via fermions is more favorable, as it naturally allows for lower temperatures and links the connection more into Standard Model particles due to their chiral nature [becker2024confronting].
In Ref. [allahverdi2010reheating], it is shown that both reheating mechanisms are rather oversimplified, as they neglect the violation of the fluctuation-dissipation theorem (see Refs. [hu1995fluctuation, berera2005absence, gleiser1994microphysical]). Additionally, these mechanisms overlook the effects of preheating and can only be significant if preheating is inefficient. This regime slowly transitions to a radiation-dominated regime, and the perturbative decay channel opens only gradually. It is implied that perturbative reheating via bosons is strongly disfavored due to the high reheating temperature, which contradicts our low reheating temperature prediction. Also, during the matter-dominated regime, all dominant particle production is mostly heavy non-relativistic [kofman1997towards]. On the other hand, fermionic reheating suffers not only from the same problem but also from the Pauli exclusion principle, which further suppresses their production. Even though this mode is more favorable than the bosonic one, its reheating scenario is still not convincing.
In addition, Ref. [risdianto2025second], which discusses inflation with a non-minimal coupling model, shows that perturbative reheating is strongly disfavored. This is because achieving a favorable reheating temperature requires violating unitarity.
@@ 7.4 Reheating temperature by preheating
As we learn from the perturbative reheating, it has some issues at certain points. Thus, the reheating temperature should be determined from the burst of particle production resulting from resonance. Our chosen mechanism is as follows: Due to the nature of the matter-dominated regime, it is expected that the reheating temperature will be evaluated at the end of the preheating stage. However, the reheating temperature should not be evaluated from the ordinary perturbative decay of either bosons or fermions, as discussed in the previous subsections. Instead, the reheating temperature should be evaluated from the last burst of scalaron oscillation that produces scalar $\chi$ at the end of the preheating stage [^7] This was discussed in Ref. [risdianto2025second]). Although in this reference, it used the beginning of the second stage of preheating.. The produced scalar $\chi$ will immediately decay into relativistic fermions $\psi$ to reheat the universe [^8] We choose the bosons $\chi$ as the product at the zero crossing of the scalaron due to the abundance of bosons over fermions during preheating. Fermions $\psi$ are chosen as the decay product of $\chi$ because they are more directly related to SM particles due to their chiral nature.. Intuitively, we can assume that during this time, the remaining energy density of the inflaton is very small. Thus, a low reheating temperature is expected.
To calculate this reheating temperature, it is written as
$T_{\text{reh}}=\left(\frac{30}{g^{*}\pi^{2}}\delta\rho\right)^{1/4},$ (7.9)
where $\delta\rho$ is the energy density produced by the single crossing during the ’last’ preheating. The $\delta\rho$ can be evaluated as [risdianto2025second, hashimoto2021inflation]
@@
Considering the chosen parameters, we need to clarify some favored values. The coupling $y$ should be small to suppress instant reheating (see [bezrukov2009initial]). For $m_{\chi}$, this lies close to the electroweak scale and should be higher. A model with this mass range favors linkage to the SM. However, as $\chi$ and fermion $\psi$ have bare masses, and their mass does not come from the symmetry breaking, they should not belong to the SM particles. [^9] As it is noted that the SM particles are massless at higher than the electroweak scale Thus, for safety, $m_{\chi}\gg m_{\psi}\gg 10^{-16}M_{p}$. In conclusion, both $\chi$ and $\psi$ masses should lie above the electroweak scale, as neither fields behave like an SM particle. Note that $y$ could be much smaller. Thus, all other parameters could be set to higher values, $k_{\chi}\gg 10^{-8}M_{p}$ and $m_{\chi}\gg 10^{-16}M_{p}$. In this model, it seems that the reheating temperature is weakly constrained by parameters $y$, $k_{\chi}$, and $m_{\chi}$.
Lastly, Table 2 presents the preferred parameter values that accommodate the constrained reheating temperature. From the table, it can be seen that a reheating temperature of approximately $10^{4}$ GeV can be achieved. These parameters can also be adjusted to yield a higher reheating temperature (e.g., $10^{8}$ GeV), which is consistent with the previous Planck results. However, when applying the bound $n_{s}=0.9743\pm 0.0034$, the expected reheating temperature decreases to only a few hundred MeV. Therefore, our proposed mechanism fails under this condition and requires further investigation in future work [Risdianto2026PreheatingStarobinsky].
@@ 8 Conclusion
In this paper, we reinvestigate Starobinsky inflation, particularly the preheating stage, following the release of the ACT results. These results suggest a significant increase in the e-fold number ($N\simeq 75$) and constrain the reheating temperature to $T_{\text{reh}}\sim 10^{4}$ GeV if we adopt the center value with $68\%$ of the confidence level. Consequently, the preheating and reheating mechanisms must be modified to align with these new results. Such a low reheating temperature affects the entire preheating stage, not just the reheating mechanism. In addition, the direct perturbative decay from the scalaron to scalar $\chi$ is strongly disfavored if the preheating mechanism is dominant and the high reheating temperature is implied. On the other hand, reheating via the decay of the scalaron to fermions $\psi$ is also disfavored if the preheating stage is dominant, although less severely, due to the possibility of achieving a lower reheating temperature and its chirality nature.
For the efficient preheating to occur, a certain initial condition of $\chi$ needs to exist at the start of preheating. This initial condition is believed to have existed because $\chi$ is a spectator field during inflation. Surprisingly, this burst of resonance particles does not occur during the inflaton zero-crossing, as many inflationary models do, but rather at the moment when tachyonic preheating takes place. However, tachyonic preheating itself does not seem to be efficient enough to drain the inflaton’s energy density at once. This is a rather peculiar condition compared with many other models of inflation.
@@
In the following, we have also considered the favored parameters in our model based on our discussion in the previous sections. In our model, $y$ should be small to avoid instant reheating, which would lead to a high reheating temperature. The non-minimal coupling $\xi$ should be treated carefully, since the preheating could not be efficient enough if $\xi\simeq 1$ and should not be greater than $10$ to avoid the unitarity issues. Also, $m_{\chi}$ should be lower than $\frac{1}{2}M$ to ensure that the scalaron decays completely and higher than the electroweak scale ($\gtrsim 10^{3}$ GeV $\approx 10^{-15}M_{p}$), as $\chi$ is shown to be different to the SM particles to preserve the viability of our model [^10] If $m_{\chi}$ is at the electroweak scale and not the Higgs field, it may have already been observed. Also, the mass is not coming from the symmetry breaking in which should not belong to SM particles in the first place.. The mass $m_{\psi}$ should be much lower than $m_{\chi}$ to ensure a successful reheating scenario [^11] As $\psi$ is supposed to reheat the universe, it should be relativistic. Thus, $m_{\psi}\ll m_{\chi}$ is preferred.. One should note that, as $\chi$ in the early stage of preheating is a heavy, non-relativistic particle, it is implied that $k_{\chi}$ should be much lower than $\xi M^{2}$. Finally, Table 3 provides the appropriate values of the parameters. It is also shown that the chosen parameters could allow both $\chi$ and $\psi$ to be detected in future colliders.
In addition, our preheating mechanism successfully describes scenarios with a lower reheating temperature but fails when the temperature drops below $10^{3}$ GeV. The uncertainty in the ACT result, $n_{s}=0.9743+0.0034$, could lead to a reheating temperature on the order of below the GeV scale (approximately $10^{2}$ MeV), under which our preheating mechanism no longer works. This issue will be addressed in future work.
@@
$\xi$ $\approx 10$ constraint: unitarity issue (upper bound), and efficient preheating (lower bound)

### v4

@@ 5 On the fermionic preheating
In this part, we will discuss the possibility that, instead of the scalar $\chi$, the scalaron is coupled with fermions. So far, we have only discussed the bosonic preheating, even though fermionic preheating is equally possible. This type of preheating has already been discussed in, e.g., Refs. [23, 22, 18, 51]. In this section, we will qualitatively discuss the possibility of this type of preheating.
If the scalaron is only coupled with certain fermions, at a large coupling [22], the preheating is similar to the bosonic case. However, in our paper, it is heavily suppressed by the Pauli exclusion principle, in contrast to the Bose-Einstein condensate (BEC) that can occur during bosonic preheating. Fermion production becomes less suppressed once the universe expands. However, even after the end of preheating, the potential of the scalaron remains quadratic ($\propto\phi^{2}$), which still favors a matter-dominated regime. Fermions are more likely to be produced at the end of the preheating phase due to the looseness of the Pauli blocking. Finally, we safely said that preheating via this mode is strongly disfavored.
@@ 6 The PBH formation from preheating
In this section, we will discuss the PBH formation due to preheating in the Starobinsky model. We will show later that the PBH formation is consistent with it being produced during the early stage of preheating.
@@
By using $\beta_{\text{PBH}}=10^{-20}$ and $\delta=\delta_{c}=0.7$ [21, 43] on Eq. (6.1), we obtain the mass variance $\sigma_{\delta}$ as
@@
For the production of PBH during preheating, the strong parametric resonance should be applied. Our assumptions implied that PBH is most likely produced when [37, 21]
@@
The power spectrum of the variable $x$ can be expressed as [38]
@@
where $k_{x}=|\textbf{k}_{x}|$ represents the comoving wavenumber associated with $x$. Note that $x_{k}$ denotes the inverse Fourier transform of $x$. The influence of preheating on $\chi$ production can be determined as [35, 37, 21]
@@
as the comoving wavenumber evaluated at the Hubble exit at the end of inflation. For $Q\gg 1$, in which it is required to get the largest production, we get $\mu_{p}=\mu_{0}\simeq\frac{1}{2\pi}\ln 3$. In addition, we will calculate the non-adiabatic power spectrum of $\chi$ as [37, 21]
@@
Note that $\vartheta$ is the angle between $p$ and $p^{\prime}$. To solve the last equation, we need to impose [21]
@@
Finally, the mass variance $\sigma_{\delta}$ during the matter-dominated era can be calculated as follows [20]
@@
Table 1: The duration of the required PBH abundance. In this table, we used $\xi=1$. Based on the table, $\Delta t$ is insensitive to the chosen parameters.
Eq. (6.15) shows the dependence on $k_{\chi}$, $\xi$, $m_{\chi}$, and $\Delta t$. In the following, we present the numerical results for Eq. (6.15) in Table 1. From our numerical analysis, by varying $k_{\chi}$ and $m_{\chi}$ with fixed $\xi$, it appears that $\Delta t$ is insensitive to the chosen parameters. The duration required to obtain a sufficient amount of PBHs is found to be $\sim 10^{7}M_{p}^{-1}$. This value is consistent with Fig. 2, which shows that the effective growth of $\chi$ production occurs after $t>10^{6}M_{p}^{-1}$. Note that in our numerical calculation, we keep $m_{\chi},k_{\chi}\ll\xi M^{2}$. In this case, the PBH constraint may not jeopardize our model. Moreover, even if $\chi$ grows larger, at some point it cannot collapse into PBHs due to the rapidly expanding universe. Thus, PBHs are suppressed beyond a certain stage [21].
@@ 7.1 The perturbative decay of the scalaron to scalar $\chi$
$\frac{\phi}{\sqrt{6}M_{p}}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi,$ (7.1)
@@
$T_{\text{reh}}\approx 3.2\times 10^{9}\text{GeV}.$ (7.4)
@@ 7.2 The perturbative decay of the scalaron to fermion $\psi$
$\mathcal{L}_{\text{Dirac}}=e^{-2\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m_{\psi})\psi,$ (7.5)
@@
$2\sqrt{\frac{2}{3}}\frac{m_{\psi}}{M_{p}}\phi\bar{\psi}\psi.$ (7.6)
@@
$\Gamma(\phi\rightarrow\bar{\psi}\psi)\simeq\frac{1}{6\pi}\frac{m_{\psi}^{2}M}{M_{p}^{2}},$ (7.7)
@@
$T_{\text{reh}}\simeq 4\times 10^{-4}m_{\psi}.$ (7.8)
By using the ACT reheating temperature constraint of $10^{4}$ GeV, $m_{\psi}\sim 10^{8}$ GeV could satisfy the prediction. Generally speaking, this mechanism is favored; however, we will see later that it is eventually disfavored.
@@ 7.3 Remarks on the perturbative mode
In our discussion about the perturbative reheating by scalar $\chi$ and fermion $\psi$, it is shown that the scalar mode is disfavored by ACT results as it produces a high reheating temperature. Furthermore, as it is enhanced by BEC, the reheating temperature from this mode could be larger than $10^{9}$ GeV. On the other hand, reheating via fermions is more favorable, as it naturally allows for lower temperatures. However, we will see shortly that both perturbative decays are disfavored in our model.
In Ref. [3], it is shown that both reheating mechanisms are rather oversimplified, as they neglect the violation of the fluctuation–dissipation theorem (see Refs. [28, 5, 19]). Additionally, these mechanisms overlook the effects of preheating and can only be significant if preheating is inefficient. In short, the mechanism considered in our paper can be summarized as follows. At the first stage of preheating, the inflaton experiences its first zero crossing and produces massive non-thermal $\chi$ particles. In our model, the production of fermions $\psi$ is suppressed by Pauli blocking; hence, immediately after inflation, the Universe is filled predominantly with massive $\chi$ particles. These massive $\chi$ particles do not decay into daughter fields, as they redshift immediately. After several oscillations, when the Universe is close to a radiation-dominated regime, the inflaton oscillation still occurs as a narrow resonance. The last burst of particle production due to the inflaton zero crossing then determines the reheating temperature. During this stage, the $\chi$ particles decay immediately into relativistic fermions $\psi$, which subsequently reheat the Universe. This assumption has been used extensively in Refs. [52, 54], and it will be discussed in more detail in Sec. 7.4.
@@ 7.4 Reheating temperature by preheating
As mentioned in the previous subsection, the reheating temperature should be evaluated by the last burst of the preheating stage before the radiation-dominated regime, and calculate the reheating mechanism based on the assumption made in 7.3.
To calculate the reheating temperature. Firstly, we need to write the relation between the energy density $\delta\rho$ produced by the single crossing during the ’last’ preheating and the reheating temperature $T_{\text{reh}}$ via
$T_{\text{reh}}=\left(\frac{30}{g^{*}\pi^{2}}\delta\rho\right)^{1/4}.$ (7.9)
Explicitly, $\delta\rho$ can be evaluated as [52, 25]
@@
Considering the chosen parameters, we need to clarify some favored values. The coupling $y$ should be small to suppress instant reheating (see [6]). For $m_{\chi}$, this lies close to the electroweak scale and should be higher. A model with this mass range favors linkage to the SM. However, as $\chi$ and fermion $\psi$ have bare masses, and their mass does not come from the symmetry breaking, they should not belong to the SM particles. [^6] As it is noted that the SM particles are massless at higher than the electroweak scale Thus, for safety, $m_{\chi}\gg m_{\psi}\gg 10^{-16}M_{p}$. In conclusion, both $\chi$ and $\psi$ masses should lie above the electroweak scale, as neither fields behave like an SM particle. Note that $y$ could be much smaller. Thus, all other parameters could be set to higher values, $k_{\chi}\gg 10^{-8}M_{p}$ and $m_{\chi}\gg 10^{-16}M_{p}$. In this model, it seems that the reheating temperature is weakly constrained by parameters $y$, $k_{\chi}$, and $m_{\chi}$.
Lastly, Table 2 presents the preferred parameter values that accommodate the constrained reheating temperature. From the table, it can be seen that a reheating temperature of approximately $10^{4}$ GeV can be achieved. These parameters can also be adjusted to yield a higher reheating temperature (e.g., $10^{8}$ GeV), which is consistent with the previous Planck results. However, when applying the bound $n_{s}=0.9743+0.0034$, the expected reheating temperature decreases to only a few hundred MeV. Therefore, our proposed mechanism fails under this condition and requires further investigation in future work [55].
@@ 8 Conclusion
In this paper, we reinvestigate Starobinsky inflation, particularly the preheating stage, following the release of the ACT results. These results suggest a significant increase in the e-fold number ($N\simeq 75$) and constrain the reheating temperature to $T_{\text{reh}}\sim 10^{4}$ GeV if we adopt the center value with $68\%$ of the confidence level. Consequently, the preheating and reheating mechanisms must be modified to align with these new results. Such a low reheating temperature affects the entire reheating mechanism related to the preheating stage. In addition, the direct perturbative decay after inflation from the scalaron to scalar $\chi$ is strongly disfavored if the preheating mechanism is dominant and the high reheating temperature is implied. On the other hand, reheating via the decay of the scalaron to fermions $\psi$ is also disfavored if the preheating stage is dominant, although less severely, due to the possibility of achieving a lower reheating temperature.
For the efficient preheating to occur, a certain initial condition of $\chi$ with $\chi(0)\gtrsim 10^{-3}M_{P}$ needs to exist at the start of preheating. This initial condition is believed to have existed because $\chi$ is a spectator field during inflation. The fluctuation of this spectator’s field during inflation is beyond the scope of this paper.
@@
In the following, we have also considered the favored parameters in our model based on our discussion in the previous sections. In our model, $y$ should be small to avoid instant reheating, which would lead to a high reheating temperature. The non-minimal coupling $\xi$ should be treated carefully, since the preheating could not be efficient enough if $\xi\simeq 1$ and should not be greater than $10$ to avoid the unitarity issues. Also, $m_{\chi}$ should be lower than $\frac{1}{2}M$ to ensure that the scalaron decays completely and higher than the electroweak scale ($\gtrsim 10^{3}$ GeV $\approx 10^{-15}M_{p}$), as $\chi$ is shown to be different to the SM particles to preserve the viability of our model [^7] If $m_{\chi}$ is at the electroweak scale and not the Higgs field, it may have already been observed. Also, the mass is not coming from the symmetry breaking, which should not belong to SM particles in the first place.. The mass $m_{\psi}$ should be much lower than $m_{\chi}$ to ensure a successful reheating scenario [^8] As $\psi$ is supposed to reheat the universe, it should be relativistic. Thus, $m_{\psi}\ll m_{\chi}$ is preferred.. One should note that, as $\chi$ is a heavy, non-relativistic particle in the early stages of preheating, $k_{\chi}$ should be much smaller than $\xi M^{2}$. Finally, Table 3 lists the appropriate parameter values. It is also shown that the chosen parameters could allow both $\chi$ and $\psi$ to be detected in future colliders.
In addition, our preheating mechanism successfully describes scenarios with a lower reheating temperature but fails when the temperature drops below $1$ GeV. The uncertainty in the ACT result, $n_{s}=0.9743+0.0034$, could lead to a reheating temperature on the order of below the GeV scale (approximately $10^{2}$ MeV), under which our preheating mechanism no longer works. This issue will be addressed in future work [55].
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the author(s) used Grammarly and ChatGPT in order to help with the grammar and clarity. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@
$\xi$ $\lesssim 10$ constraint: unitarity issue (upper bound), and efficient preheating (lower bound)